# Jean DELHAYE

### v2

@@ 3.3. Absolute continuity
In Subsection 3.1, we have shown that any Brownian motion on $U_{N}^{+}$ coming from a multiple of the Laplace-Beltrami operator of $U_{N}$ has its constants $\alpha$ and $\beta$ linked by:
$\beta\sim\frac{\alpha}{N},\quad\hbox{as }N\longrightarrow\infty.$
In consequence, we will only consider Brownian motions where $\beta=O(\alpha/N)$. Let $(\psi_{t})_{t\geq 0}$ be a Lévy process associated to such a generating functional, i.e. we have
$\psi_{t}:\chi^{\epsilon}_{\mathbf{n}}\longmapsto d_{\mathbf{n}}\exp\left(-t\left(\left(\alpha-\frac{\beta}{N}\right)\lambda_{\mathbf{n}}+\beta\frac{\mathfrak{p}_{\mathbf{n}}+2\mathfrak{e}_{\mathbf{n}}}{N^{2}}\right)\right),\quad t\geq 0,$
for some $\alpha,\beta\geq 0$ with $\beta=O(\alpha/N)$, where we have written $\lambda_{\mathbf{n}}=P^{\prime}_{\mathbf{n}}(N)/P_{\mathbf{n}}(N)$.
In this paper, to compute limit profiles, we only consider the total variation distance. Just like in the classical setting, the total variation distance for states is much easier to compute when one of the state is absolutely continuous w.r.t the other. A state $\psi$ is said to be absolutely continuous w.r.t the Haar state if there exists $f\in\mathrm{L}^{1}(U_{N}^{+})$ such that $\psi(x)=h(fx)$ for all $x\in\mathcal{O}(U_{N}^{+})$. In this case, the total variation distance is easily expressed (see [21, Lem 2.6])
@@
Quantum groups exhibit a unique characteristic wherein absolute continuity is not easily obtained. Let us first present an elementary fact before stating a proposition on absolute continuity.
@@
$0\leq 2\mathfrak{e}_{\mathbf{n}}+\mathfrak{p}_{\mathbf{n}}\leq\frac{\ell(\mathbf{n})^{2}}{4}.$
@@
We will use the notations introduced immediately before Lemma 3.5. Let us set $m_{i}:=k_{i}+i$ ($1\leq i\leq\mathfrak{p}_{\mathbf{n}}$). We further denote by $x$ and $y$ the amount of even and odd numbers among the $m_{i}$’s respectively. We thus have
$2\mathfrak{e}_{\mathbf{n}}+\mathfrak{p}_{\mathbf{n}}$ $=2\sum_{1\leq i<j\leq\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{j}+j-(k_{i}+i)}+x+y$
$=2\left(\frac{x(x-1)+y(y-1)}{2}-xy\right)+x+y$
@@
this shows the positivity. To show the upper bound, it suffices to notice that if $m_{j}$ and $m_{j+1}$ have the same parity, then we have $k_{j+1}>k_{j}+1$. This further implies that there are at most $\lfloor\ell(\mathbf{n})/2\rfloor$ odd (or even) numbers among the $m_{j}$’s. This allows us to conclude
$2\mathfrak{e}_{\mathbf{n}}+\mathfrak{p}_{\mathbf{n}}=(x-y)^{2}$ $\leq\max(x,y)^{2}\leq\frac{\ell(\mathbf{n})^{2}}{4}.$
@@
Fix $c>0$ and $t=\alpha^{-1}(N\ln(\sqrt{2}N)+cN)$. Provided $N$ is large enough, then $\psi_{t}$ is absolutely continuous w.r.t. the Haar state.
@@
We want to see that the series
@@
$f_{t}:=\sum_{\epsilon,\mathbf{n}}d_{\mathbf{n}}\exp\left(-t\left(\left(\alpha-\frac{\beta}{N}\right)\lambda_{\mathbf{n}}+\beta\frac{\mathfrak{p}_{\mathbf{n}}+2\mathfrak{e}_{\mathbf{n}}}{N^{2}}\right)\right)\chi_{\mathbf{n}}^{\epsilon}$
@@
converges in $\mathrm{L}^{1}(U_{N}^{+})$ for $N$ large enough. We will make use of the $\mathrm{L}^{2}$-norm, as the $\mathrm{L}^{1}$-norm is dominated by it. Take $N$ large enough so that
$\frac{\beta t}{N^{2}}\leq\frac{c}{2}.$
@@
Using the estimates from [17, Lem 1.7], Lemma 3.7 and the fact that $d_{n}\leq N^{n}$, we have
$\|f_{t}\|_{2}^{2}\leq\sum_{\epsilon,\mathbf{n}}d_{\mathbf{n}}^{2}e^{-2t(\alpha-\beta/N)\lambda_{\mathbf{n}}}\leq 2\sum_{m\geq 1}\sum_{\begin{subarray}{c}\mathbf{n}=(n_{1},...,n_{p}),\\ n_{1}+...+n_{p}=m\end{subarray}}N^{2m}e^{-2t(\alpha-\beta/N)m/N}\leq\sum_{m\geq 1}e^{-cm}<\infty.$
4. Restricting the study to a smaller algebra
Let us recall that the asymptotic study of the Brownian motion on the orthogonal quantum group $O_{N}^{+}$ was done in [23]. However, the proof heavily relies on the commutativity of the central algebra $\mathcal{O}(O_{N}^{+})_{0}$, this is not the case for $\mathcal{O}(U_{N}^{+})_{0}$ (note for instance that $\chi_{1}^{2}\neq\overline{z}\chi_{1}^{2}z$). Our first goal will be to find a smaller commutative subalgebra on which we may restrict our study to drastically ease the computations. We proceed by understanding what would be the smallest algebra $A\subset\mathcal{O}(U_{N}^{+})$ on which all the information of the limit-profile is contained. Before explicitly introducing this algebra, let us present some facts.
@@ 4. Restricting the study to a smaller algebra
For any tuple $\mathbf{n}$, we have, setting $t_{c}=N\ln(\sqrt{2}N)+cN$ for some $c\in\mathbb{R}$:
$\psi_{t_{c}}(\chi_{\mathbf{n}}^{\epsilon})\underset{N\to\infty}{\longrightarrow}e^{-\widetilde{c}|\mathbf{n}|},\quad\hbox{with }\widetilde{c}=c+\ln\sqrt{2}$
@@
This follows from the fact that $d_{n_{j}}e^{-t_{c}\lambda_{n_{j}}}\to e^{-\widetilde{c}n_{j}}$ as $N\to\infty$ for any $j\in\{1,...,p\}$ (see the proof of [23, Prop 3.10]). ∎
Now, we understand that the limit-process evaluated on an irreducible character $\chi_{\mathbf{n}}^{\epsilon}$ only depends on $|\mathbf{n}|$, let us introduce compositions. A composition of $m$ is a tuple $\mathbf{n}=(n_{1},...,n_{p})$ that satisfies $|\mathbf{n}|:=n_{1}+...+n_{p}=m$.
$\Pi_{m}^{1}=\{\mathbf{n}\in\Pi_{m}:n_{1}=1\}\quad\hbox{\&}\quad\Pi_{m}^{>0}=\Pi_{m}\setminus\Pi_{m}^{1}.$
One easily checks that the following maps
$\left\{\begin{matrix}\Pi_{m}&\longrightarrow&\Pi_{m+1}^{1}\\ \mathbf{n}&\longmapsto&(1,\mathbf{n})\end{matrix}\right.,\quad\left\{\begin{matrix}\Pi_{m}&\longrightarrow&\Pi_{m+1}^{>0}\\ \mathbf{n}&\longmapsto&\mathbf{n}^{+}\end{matrix}\right.,$
$\left\{\begin{matrix}\Pi_{m+1}^{1}&\longrightarrow&\Pi_{m}\\ \mathbf{n}&\longmapsto&\mathbf{n}^{-}\end{matrix}\right.\quad\hbox{\&}\quad\left\{\begin{matrix}\Pi_{m+1}^{>0}&\longrightarrow&\Pi_{m}\\ \mathbf{n}&\longmapsto&\mathbf{n}^{-}\end{matrix}\right.,$
are bijections (where we have written $\mathbf{n}^{\pm}=(n_{1}\pm 1,n_{2},...,n_{p})$ and identified $(0,\mathbf{n})\sim\mathbf{n}$).
$x:=\frac{\overline{z}\chi_{1}+\chi_{1}z}{\sqrt{2}},$
let us denote it by $\mathcal{O}(U_{N}^{+})_{00}$. First, the restriction of the Haar state coincides with integration w.r.t. the semicircle distribution $\nu_{\mathrm{SC}}$ (see [3, Thm 1]). More precisely, we have $h=\nu_{\mathrm{SC}}\circ\iota$ where $\iota:\mathcal{O}(U_{N}^{+})_{00}\to\mathbb{C}[X]$ is the isomorphism given by $P_{m}(x)=:x_{m}\mapsto P_{m}$. Before stating and proving this section’s final result, let us prove a quick but useful lemma.
Lemma 4.2.
Let $\mathbf{n}$ be any tuple of integer and $\epsilon\in\{\pm 1\}$, we have
$x\cdot\chi_{\mathbf{n}}^{\epsilon}=\frac{\chi_{(1,\mathbf{n})}^{\epsilon}+\chi_{\mathbf{n}^{+}}^{-\epsilon}+\chi_{\mathbf{n}^{-}}^{-\epsilon}}{\sqrt{2}},$

### v3

@@ 7.1. Absolute continuity
To compute limit profiles, we only consider the total variation distance. Just like in the classical setting, the total variation distance for states is much easier to compute when one of the states is absolutely continuous w.r.t. the other. A state $\psi$ is said to be absolutely continuous w.r.t. the Haar state if there exists $f\in\mathrm{L}^{1}(U_{N}^{+})$ such that $\psi(x)=h(fx)$ for all $x\in\mathcal{O}(U_{N}^{+})$. In this case, the total variation distance is easily expressed (see [Fre19, Lem 2.6])
@@
Quantum groups exhibit a unique characteristic where absolute continuity is not easily obtained. Let us first present an elementary fact before stating a proposition on absolute continuity.
@@
$0\leq\mathfrak{e}_{\mathbf{n}}\leq\ell(\mathbf{n})^{2}.$
Moreover, the upper bound is attained if and only if the tuple consists only of odd numbers.
@@
We will use the notations introduced immediately before Lemma 3.6. Let us set $x/y$ to be the amount of even/odd numbers within the set $\{k_{i}+i:1\leq i\leq\mathfrak{p}_{\mathbf{n}}\}$. We thus have
$\mathfrak{e}_{\mathbf{n}}$ $=\sum_{1\leq i,j\leq\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{j}+j-(k_{i}+i)}$
$=\left(\sum_{i=1}^{\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{i}+i}\right)^{2}$
@@
this shows the positivity. The upper bound is clear:
$\mathfrak{e}_{\mathbf{n}}=(x-y)^{2}$ $\leq\max(x,y)^{2}\leq\mathfrak{p}_{\mathbf{n}}^{2}\leq\ell(\mathbf{n})^{2}.$
It is clearly attained when $\mathbf{n}$ consists only of odd numbers, as this would imply $k_{i}+i=2i$ for all $1\leq i\leq\mathfrak{p}_{\mathbf{n}}$, leading to $x=\mathfrak{p}_{\mathbf{n}}=\ell(\mathbf{n})$ and $y=0$. If $\mathbf{n}$ has at least one even entry, simply observe that
$\max(x,y)\leq\mathfrak{p}_{\mathbf{n}}<\ell(\mathbf{n}).$
@@
Let $c\in\mathbb{R}$ and set $t_{N}=\alpha^{-1}(N\ln(\sqrt{2}N)+cN)$, then
  • if $c>0$, then $\psi_{t_{N}}$ is absolutely continuous w.r.t. the Haar state, provided $N$ is large enough;
  • if $c<0$, then $\psi_{t_{N}}$ is not absolutely continuous w.r.t. the Haar state, provided $N$ is large enough.
@@
Assume $c>0$. We want to see that the series
@@
$f_{t_{N}}:=1+\sum_{\epsilon,\mathbf{n}}d_{\mathbf{n}}\exp\left(-t_{N}\left(\alpha\lambda_{\mathbf{n}}+\beta\frac{\mathfrak{e}_{\mathbf{n}}}{N}\right)\right)\chi_{\mathbf{n}}^{\epsilon}$
@@
converges in $\mathrm{L}^{1}(U_{N}^{+})$ for $N$ large enough. We will make use of the $\mathrm{L}^{2}$-norm, as the $\mathrm{L}^{1}$-norm is dominated by it. Using the estimates from [FHLUZ17, Lem 1.7], Lemma 7.1 and the fact that $d_{n}\leq N^{n}$, we have
$\|f_{t}-1\|_{2}^{2}$ $\leq\sum_{\epsilon,\mathbf{n}}d_{\mathbf{n}}^{2}e^{-2t_{N}\alpha\lambda_{\mathbf{n}}}$
$\leq 2\sum_{m\geq 1}\sum_{\mathbf{n}\in\Pi_{m}}N^{2m}e^{-2t_{N}\alpha m/N}$
$\leq 2\sum_{m\geq 1}|\Pi_{m}|e^{-2m(c+\ln\sqrt{2})}$
$\leq\sum_{m\geq 1}e^{-2mc}$
$<\infty.$
For the case $c<0$, we restrict to the commutative subalgebra $\mathcal{O}(U_{N}^{+})_{00}$ (see Proposition 4.3), we may view the state $\psi_{t_{N}}$ and the Haar state as probability measures $\mu_{t_{N}}$ and $\nu_{\mathrm{SC}}$ respectively via the isomorphism $\mathcal{O}(H_{N}^{s+})_{00}\to\mathbb{C}[X],x\mapsto X$. Propositions 4.3 and 5.3 then imply
$\mu_{t_{N}}(P_{2k})$ $=\psi_{t_{N}}(x_{2k})$
$=\frac{2}{2^{k}}\sum_{\mathbf{n}\in\Pi_{2k}}\psi_{t_{N}}(\chi_{\mathbf{n}}^{\epsilon})$
$\geq\frac{2}{2^{k}}\sum_{\mathbf{n}\in A_{2k}^{0}}d_{\mathbf{n}}e^{-t_{N}\alpha\lambda_{\mathbf{n}}}$
$\underset{N\to\infty}{\longrightarrow}\frac{1}{2^{k}}\binom{2k}{k}e^{-2k(c+\ln\sqrt{2})}$
$=\frac{e^{-2kc}}{2^{2k}}\binom{2k}{k}$
$=\eta_{c}^{\infty}(P_{2k}).$
This further implies that
$\underset{N\to\infty}{\lim\inf}\,\,\mu_{t_{N}}(X^{2n})\geq\eta_{c}^{\infty}(X^{2n}),\quad n\in\mathbb{N}.$
This is because $X^{2n}$ can be expressed as a linear combination of $P_{2k}$’s with positive coefficients. Now, since $\eta_{c}^{\infty}$ allocates mass outside of [2,2][-2,2] (see Proposition 5.3), then so must $\mu_{t_{N}}$ for $N$ large enough. Recall that through the considered isomorphism, the Haar state corresponds to the semicircle distribution whose support is [2,2][-2,2], hence absolute continuity does not hold. ∎
7.2. Computation of the limit profile
$\psi_{t}(\chi_{2})=(N^{2}-1)e^{-2tN\alpha/(N^{2}-1)}\neq N^{2}e^{-2t(N\alpha+2\beta)/N^{2}}=\psi_{t}(\chi_{(1,1)}).$
We will make use of the process defined by $\widetilde{\psi}_{t}=\psi_{t}\circ\mathbb{F}$.
Theorem 7.3.
Let $\psi^{(\alpha,\beta)}=(\psi_{t})_{t\geq 0}$ be a Brownian motion of parameters $(\alpha,\beta)\in\mathbb{R}_{+}^{2}$ on $U_{N}^{+}$, assume that $\alpha>0$ and that
$\frac{\beta\ln N}{\alpha}\underset{N\to\infty}{\longrightarrow}r,$
for some $r\in[0,\infty]$. Then, $\psi^{(\alpha,\beta)}$ has cutoff at time $\alpha^{-1}N\ln N$. More precisely, it exhibits the following cutoff:
$d_{\mathrm{TV}}\big(\widetilde{\psi}_{\alpha^{-1}(\ln(\sqrt{2}N)+cN},h\big)\underset{N\to\infty}{\longrightarrow}d_{\mathrm{T}V}(\eta_{c}^{r},\nu_{\mathrm{SC}}),\quad c>0$
and
$\underset{N\to\infty}{\lim\inf}\,d_{\mathrm{TV}}\big(\widetilde{\psi}_{\alpha^{-1}(\ln(\sqrt{2}N)+cN},h\big)\geq d_{\mathrm{T}V}(\eta_{c}^{r},\nu_{\mathrm{SC}}),\quad c\in\mathbb{R},$
where $\nu_{\mathrm{SC}}$ is the semicircle distribution and $\eta_{c}^{r}$ is the probability measure introduced in Lemma 5.1.
Proof.
Up to renormalising, we may assume that $\alpha=1$. We first prove the result for the process $(\widetilde{\psi}_{t})_{t\geq 0}=(\psi_{t}\circ\mathbb{F})_{t\geq 0}$. The process being $\mathbb{F}$-invariant, we may restrict the study to the commutative algebra $\mathcal{O}(U_{N}^{+})_{00}$ for which the $x_{m}$’s form an orthonormal basis.
Let $c>0$ and set $t_{N}=N\ln(\sqrt{2}N)+cN$. It follows from Proposition 7.2 that $\widetilde{\psi}_{t_{N}}$ has an $\mathrm{L}^{1}$-density (at least asymptotically defined) given by
$\widetilde{f}_{t_{N}}=\sum_{m\geq 0}{\widetilde{\psi}_{t_{N}}(x_{m})}x_{m}=1+\sum_{m\geq 1}\frac{1}{\sqrt{2}^{m}}\left(\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}\psi_{t_{N}}(\chi^{\epsilon}_{\mathbf{n}})\right)x_{m}.$
Using the isomorphism $\mathcal{O}(U_{N}^{+})_{00}\to\mathbb{C}[X],x_{m}\mapsto P_{m}$ sending the Haar state to the semicircle distribution, we have
$d_{\mathrm{TV}}(\widetilde{\psi}_{t_{N}},h)=\frac{1}{2}\|\widetilde{f}_{t_{N}}-1\|_{1,h}=\frac{1}{2}\left\|\sum_{m\geq 1}\widetilde{\psi}_{t_{N}}(x_{m})P_{m}\right\|_{1,\nu_{\mathrm{SC}}}.$
In addition, Proposition 5.3 implies that:
$\widetilde{\psi}_{t_{N}}(x_{m})\underset{N\to\infty}{\longrightarrow}e^{-cm}\mathbb{E}[e^{-rS_{n}^{2}}],\quad m\in\mathbb{N},$
@@ 7.2. Computation of the limit profile
where $(S_{n})_{n\geq 0}$ is a $\pm 1$ walk. Moreover, for any fixed $m\in\mathbb{N}$, we have using yet again the estimates from [FHLUZ17, Lem 1.7], Lemma 7.1 and the fact that $d_{n}\leq N^{n}$,
$\|\widetilde{\psi}_{t_{N}}(x_{m})x_{m}\|_{1}$ $\leq\|\widetilde{\psi}_{t_{N}}(x_{m})x_{m}\|_{2}$
$=\frac{1}{\sqrt{2}^{m}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}\psi_{s_{t}}(\chi^{\epsilon}_{\mathbf{n}})$
$\leq\frac{1}{\sqrt{2}^{m}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}d_{\mathbf{n}}e^{-t_{N}\lambda_{\mathbf{n}}}$
$\leq\sqrt{2}^{m}N^{m}e^{-m(\ln(\sqrt{2}N)+c)}$
$=e^{-cm}.$
The bound being uniform in $N$ allows for an exchange of the sum over $m$ and of the limit in $N$, yielding using Lemma 5.1 and Proposition 5.3:
$\lim_{N\to\infty}d_{\mathrm{TV}}(\widetilde{\psi}_{t_{N}},h)$ $=\frac{1}{2}\left\|\sum_{m\geq 1}e^{-cm}\mathbb{E}[e^{-rS_{n}^{2}}]P_{m}\right\|_{1,\nu_{\mathrm{sc}}}$
$=\frac{1}{2}\left\|\sum_{m\geq 1}\eta_{c}^{r}(P_{m})P_{m}\right\|_{1,\nu_{\mathrm{sc}}}$
$=d_{\mathrm{TV}}(\eta_{c}^{r},\nu_{\mathrm{SC}})$
Now, for the second result, fix $c\in\mathbb{R}$, set $t_{N}=N\ln(\sqrt{2}N)+cN$ and consider the states $\widetilde{\psi}_{t_{N}}$ and $h$ in their measure form as $\mu_{t_{N}}$ and $\nu_{\mathrm{SC}}$. As already mentioned, we have
$\mu_{t_{N}}(P_{m})\leq e^{-cm}=\eta_{c}^{0}(P_{m}),\quad m\in\mathbb{N}.$
In particular, this implies that:
$\mu_{t_{N}}(X^{2n})\leq\eta^{0}_{c}(X^{2n}),\quad n\in\mathbb{N}.$
This follows from the fact that $X^{2n}$ can be expressed as a linear combination of $P_{m}$’s with positive coefficients. Since $\eta_{c}^{0}$ is supported on the interval $\gamma_{c},\gamma_{c}]$, where $\gamma_{c}=e^{c}+e^{-c}$, it follows that $\mu_{t_{N}}$ must also be supported on this interval (as the support of a measure can be estimated by bounding its moments. See for example the beginning of [Sch20, Lect 2]). Given that moment convergence with compact support implies weak convergence, we have:
$\mu_{t_{N}}(B)\underset{N\to\infty}{\longrightarrow}\eta_{c}^{r}(B),$
for any continuity set $B$ of $\eta_{c}$, i.e., a set such that $\eta_{c}(\partial B)=0$. Since $f_{c}^{r}$ is continuous, the set $B=\{f_{c}^{r}<1\}$ is open and maximizes the total variation distance, leading to
$d_{\mathrm{TV}}(\widetilde{\psi}_{t_{N}},h)$ $\geq d_{\mathrm{TV}}(\mu_{t_{N}},\nu_{\mathrm{SC}})$
$\geq\nu_{\mathrm{SC}}(B)-\mu_{t_{N}}(B)$
$\underset{N\to\infty}{\longrightarrow}\nu_{\mathrm{SC}}(B)-\eta_{c}^{r}(B)$
$=d_{\mathrm{TV}}(\nu_{\mathrm{SC}},\eta_{c}^{r}).$
Now we make the link with the actual Brownian motion $\psi^{(\alpha,\beta)}$.
For the first convergence (setting $t_{N}=N\ln(\sqrt{2}N)+cN$ for fixed $c>0$) it suffices to see that $d_{\mathrm{TV}}\big(\psi_{t_{N}},\widetilde{\psi}_{t_{N}}\big)\to 0$ as $N\to\infty$. This is obvious as
$d_{\mathrm{TV}}\big(\psi_{t_{N}},\widetilde{\psi}_{t_{N}}\big)\leq\sum_{\epsilon,\mathbf{n}}|\psi_{t_{N}}(\chi_{\mathbf{n}}^{\epsilon})-\widetilde{\psi}_{t_{N}}(\chi_{\mathbf{n}}^{\epsilon})|^{2}\leq\|f_{t_{N}}-1\|_{2}^{2}+\|\widetilde{f}_{t_{N}}-1\|_{2}^{2}\leq 2\sum_{m\geq 1}e^{-2cm}$
where the last inequality holds for $N$ large enough (similarly to the proof of Proposition 7.2). By dominated convergence, we may exchange summation and limit in $N$ yielding the convergence as $\psi^{(\alpha,\beta)}$ and $\widetilde{\psi}$ have same moment convergence.
For the second, simply note that
$d_{\mathrm{TV}}(\psi_{t},h)=\frac{1}{2}\|\psi_{t_{c}}-h\|_{FS}\geq\frac{1}{2}\|\mathbb{F}\circ\psi_{t}-\mathbb{F}\circ h\|_{FS}=d_{\mathrm{TV}}(\widetilde{\psi}_{t},h).$
Let us make a few comments on the parameter $r=\lim_{N}\beta\ln N/\alpha$. First, note that even if no such convergence occurs, then one still has cutoff at the same time and now different subsequences may have different cutoff profiles. However, we emphasize that it is quite natural to consider Brownian motions where such a parameter exists.
Second, recall that (see Remark 7.5) cutoff profiles are only unique up to affine transformation. The end of this subsection is dedicated to proving that every $r\in[0,\infty]$ does define a unique cutoff profile.
We first estimate the weight of the singular part of the measure $\eta_{c}^{r}$.
@@
For every $r\in(0,\infty]$ set $R_{r}=\cos T_{r}$ where $T_{r}$ is a random variable of Gaussian distribution $\mathcal{N}(0,2r)$ if $r\in(0,\infty)$, and $T_{\infty}\sim\mathrm{Unif}([0,2\pi])$. Then, for every $r\in(0,\infty)$, we may write
$\mathbb{E}\left[\left(1-\epsilon^{2}R_{r}^{-2}\right)_{+}\right]=1+A_{r}\epsilon+B_{r}\epsilon^{2}+o(\epsilon^{2}),\quad\epsilon\downarrow 0,$
where
$A_{r}=-\frac{4}{\pi}\frac{1-e^{-4r^{2}}}{1+e^{-4r^{2}}}\emph{}B_{r}=\frac{4e^{-4r^{2}}}{(1+e^{-4r^{2}})^{2}}.$
Moreover,
$\mathbb{E}\left[\left(1-\epsilon^{2}R_{r}^{-2}\right)_{+}\right]=\frac{2\left(\arccos\epsilon-\epsilon\sqrt{1-\epsilon^{2}}\right)}{\pi}=1-\frac{4\epsilon}{\pi}+\frac{\epsilon^{3}}{3\pi}+O(\epsilon^{5}).$
We prove that every parameter $r\in[0,\infty]$ defines a unique cutoff profile by using the weight of the singular part.
@@
Fix $r\in(0,\infty)$ and set [^9] This is the Poisson summation formula for the heat kernel, see [Woi20].
$\phi_{r}(\theta)=\frac{1}{\sqrt{4r\pi}}\sum_{k\in\mathbb{Z}}e^{-(\theta+k\pi)^{2}/4r}=\frac{1}{\pi}\left(1+2\sum_{n\geq 1}e^{-4r^{2}n}\cos(2n\theta)\right).$
Note the following properties of the function $\phi_{r}$ which we will repetitively use through the rest of the proof:
  • $\phi_{r}$ is smooth;
  • $\phi_{r}$ is even and $\pi$-periodic;
  • $\phi_{r}(\theta)\mathbf{1}_{|\theta|<\pi/2}\mathrm{d}\theta$ defines a probability measure.
Now compute with $\epsilon>0$:
$\mathbb{E}\left[\left(1-\epsilon^{2}R_{r}^{-2}\right)_{+}\right]$ $=\frac{1}{\sqrt{4\pi r}}\int_{\{|\cos t|>\epsilon\}}\left(1-\frac{\epsilon^{2}}{\cos(t)^{2}}\right)e^{-t^{2}/4r}\mathrm{d}t$
$=\int_{-\pi/2}^{\pi/2}\left(1-\frac{\epsilon^{2}}{\cos(\theta)^{2}}\right)\phi_{r}(\theta)\mathbf{1}_{\{|\cos t|>\epsilon\}}\mathrm{d}\theta$
$=\int_{-\arccos\epsilon}^{\arccos\epsilon}\left(1-\frac{\epsilon^{2}}{\cos(\theta)^{2}}\right)\phi_{r}(\theta)\mathrm{d}\theta$
$=I_{r}^{\epsilon}-\epsilon^{2}(J_{r}^{\epsilon}+K_{r}^{\epsilon}),$
with
$I_{r}^{\epsilon}$ $=\int_{-\arccos\epsilon}^{\arccos\epsilon}\phi_{r}(\theta)\mathrm{d}\theta$
$=1-2\int_{\arccos\epsilon}^{\pi/2}\phi_{r}(\theta)\mathrm{d}\theta$
$=1-2\int_{0}^{\arcsin\epsilon}\phi_{r}\left(\frac{\pi}{2}-\theta\right)\mathrm{d}\theta$
$=1-2\int_{0}^{\arcsin\epsilon}\phi_{r}\left(\frac{\pi}{2}+\theta\right)\mathrm{d}\theta$
$=1-2\left(\phi_{r}\left(\frac{\pi}{2}\right)\arcsin\epsilon+\frac{\phi_{r}^{\prime}(\pi/2)}{2}\arcsin(\epsilon)^{2}+O\left(\arcsin(\epsilon)^{3}\right)\right)$
$=1-2\phi_{r}\left(\frac{\pi}{2}\right)\epsilon+O(\epsilon^{3}).$
From the third to fourth line we have used the fact that $\phi_{r}^{\prime}(\pi/2)=0$. Now for the two last terms
$J_{r}^{\epsilon}$ $=\int_{-\arccos\epsilon}^{\arccos\epsilon}\frac{\phi_{r}(\pi/2)}{\cos(\theta)^{2}}\mathrm{d}\theta$
$=2\phi_{r}\left(\frac{\pi}{2}\right)\tan\big(\arccos\epsilon\big)$
$=2\phi_{r}\left(\frac{\pi}{2}\right)\frac{\sqrt{1-\epsilon^{2}}}{\epsilon}$
$=\frac{2}{\epsilon}\phi_{r}\left(\frac{\pi}{2}\right)+O(\epsilon),$
and
$K_{r}^{\epsilon}$ $=\int_{-\arccos\epsilon}^{\arccos\epsilon}\frac{\phi_{r}(\theta)-\phi_{r}(\pi/2)}{\cos(\theta)^{2}}\mathrm{d}\theta$
$=2\int_{0}^{\arccos\epsilon}\frac{\phi_{r}(\theta)-\phi_{r}(\pi/2)}{\cos(\theta)^{2}}\mathrm{d}\theta$
$=2\int_{0}^{\pi/2}\frac{\phi_{r}(\theta)-\phi_{r}(\pi/2)}{\cos(\theta)^{2}}\mathrm{d}\theta-2\int_{\arccos\epsilon}^{\pi/2}\frac{\phi_{r}(\theta)-\phi_{r}(\pi/2)}{\cos(\theta)^{2}}\mathrm{d}\theta$
$=\frac{4}{\pi}\sum_{n\geq 1}e^{-4r^{2}n}\int_{0}^{\pi/2}\frac{\cos(2n\theta)-(-1)^{n}}{\cos(\theta)^{2}}\mathrm{d}\theta+O(\epsilon)$
$=\frac{4}{\pi}\sum_{n\geq 1}e^{-4r^{2}n}\int_{0}^{\pi/2}\frac{\cos(2n\theta)-(-1)^{n}}{\cos(\theta)^{2}}\mathrm{d}\theta+O(\epsilon)$
$=4\sum_{n\geq 1}ne^{-4r^{2}n}(-1)^{n}+O(\epsilon).$
From the fifth to sixth line we have computed the following integral:
$\int_{0}^{\pi/2}\frac{\cos(2n\theta)-(-1)^{n}}{\cos(\theta)^{2}}\mathrm{d}\theta$ $=\int_{0}^{\pi/2}\frac{\cos(2n(\pi/2-\theta))-(-1)^{n}}{\cos(\pi/2-\theta)^{2}}\mathrm{d}\theta$
$=\int_{0}^{\pi/2}\frac{(-1)^{n}\cos(2n\theta)-(-1)^{n}}{\sin(\theta)^{2}}\mathrm{d}\theta$
$=(-1)^{n}\int_{0}^{\pi/2}\frac{\cos(2n\theta)-1}{\sin(\theta)^{2}}\mathrm{d}\theta$
$=(-1)^{n+1}\int_{0}^{\pi/2}\frac{2\sin(n\theta)^{2}}{\sin(\theta)^{2}}\mathrm{d}\theta$
$=(-1)^{n+1}\int_{-\pi/2}^{\pi/2}\left(\frac{e^{\mathrm{i}n\theta}-e^{-\mathrm{i}n\theta}}{e^{\mathrm{i}\theta}-e^{-\mathrm{i}\theta}}\right)^{2}\mathrm{d}\theta$
$=(-1)^{n+1}\int_{-\pi/2}^{\pi/2}e^{2\mathrm{i}(n-1)\theta}\left(\frac{1-e^{-2\mathrm{i}n\theta}}{1-e^{-2\mathrm{i}\theta}}\right)^{2}\mathrm{d}\theta$
$=(-1)^{n+1}\int_{-\pi/2}^{\pi/2}e^{2\mathrm{i}(n-1)\theta}\sum_{k,\ell=0}^{n-1}e^{-2\mathrm{i}(k+\ell)\theta}\mathrm{d}\theta$
$=(-1)^{n+1}\pi n.$
Putting everything together gives the result for $r$ finite:
$A_{r}$ $=-4\phi_{r}\left(\frac{\pi}{2}\right)$
$=-\frac{4}{\pi}\left(1+2\sum_{n\geq 1}(-1)^{n}e^{-4r^{2}n}\right)$
$=-\frac{4}{\pi}\left(1-2\frac{e^{-4r^{2}}}{1+e^{-4r^{2}}}\right)$
$=-\frac{4}{\pi}\frac{1-e^{-4r^{2}}}{1+e^{-4r^{2}}},$
and
$B_{r}$ $=-4\sum_{n\geq 1}ne^{-4r^{2}n}(-1)^{n}$
$=\frac{4e^{-4r^{2}}}{(1+e^{-4r^{2}})^{2}}.$
We now treat the infinite case.
$\mathbb{E}\left[\left(1-\epsilon^{2}R_{\infty}^{-2}\right)_{+}\right]$ $=\frac{1}{\pi}\int_{-\pi/2}^{\pi/2}\left(1-\frac{\epsilon^{2}}{\cos(\theta)^{2}}\right)_{+}\mathrm{d}\theta$
$=\frac{1}{\pi}\int_{-\arccos\epsilon}^{\arccos\epsilon}\left(1-\frac{\epsilon^{2}}{\cos(\theta)^{2}}\right)\mathrm{d}\theta$
$=\frac{2}{\pi}\left(\arccos\epsilon-\epsilon^{2}\tan(\arccos\epsilon)\right)$
$=\frac{2}{\pi}\left(\arccos\epsilon-\epsilon\sqrt{1-\epsilon^{2}}\right)$
$=1-\frac{4\epsilon}{\pi}+\frac{\epsilon^{3}}{3\pi}+O(\epsilon^{5}).$
Proposition 7.5.
Every $r\in[0,\infty]$ defines a unique cutoff profile through the mapping
$g_{r}:c\longmapsto d_{\mathrm{TV}}(\nu_{\mathrm{SC}},\eta_{c}^{r}).$